# A new characterization of (pre)liminary C*-algebras

### v3

@@ 3.2. Hausdorff decompositions
We introduce the notion of standard topological space. Recall the notion of standard Borel space from [34, Section 12.B].
@@
This follows from [34, Lemma 13.2 and Lemma 13.3] considering that $X$ has a countable basis of open sets. ∎
@@
This follows from [34, Theorem 6.9]. ∎
@@ 3.3. Constructible sets and Baire spaces
We consider the following terminology about collections of subsets of a given set; see [29, Section 4].
@@
Let $X$ be a topological space. A subspace $A$ of $X$ is called meager if it is contained in the union of a countable family of closed sets with empty interior. A topological space $X$ is Baire if every nonempty open subset $U$ of $X$ is not meager; see [34, Definition 8.2]. A topological space $X$ is hereditarily Baire if each of its closed subspace of $X$ is Baire, when endowed with the subspace topology. It is easily seen by induction on the complexity of $C$ that, if $X$ is a hereditarily Baire space, and $C\subseteq X$ is constructible, then $C$ is hereditarily Baire.
@@ 3.4. Polish presentations
A space with a standard Polish presentation is a space $X$ explicitly presented as the quotient $P/E$ of a Polish space $P$ by an equivalence relation $E$ with the property that:
@@
    the quotient map $P\rightarrow P/E$ is (continuous and) open when $X$ is endowed with the quotient topology, i.e., the $E$-saturation of an open subset of $P$ is open;
@@
    the quotient Borel structure on $P/E$ induced by the Borel structure on $P$ is standard, and in particular $E$ is a smooth equivalence relation [34, Exercise 18.20].
A space with a standard Polish presentation has a countable basis of open sets, and it is hereditarily Baire. This follows from the fact that every Polish space is second countable and a Baire space [34, Theorem 8.4], and every $G_{\delta}$ subspace of a Polish space is Polish [34, Theorem 3.11].
@@ 3.5. Countably compactly based spaces
We consider the following property for topological spaces, which simultaneously subsumes second coutability and local compactness, and it is equivalent to their conjunction in the Hausdorff case. For a set $Y\subseteq X$ we denote by $\mathrm{int}\left(Y\right)$ its interior.
@@
We say that $X$ has a countable basis of compact open set when one can take the elements of $\mathcal{B}$ both compact and open.
@@ 4.1. Fell spaces and their compactification
where $K\subseteq X$ is compact and $U$ is open; see also [17, 3.9.2]. In the Hausdorff case, this is also considered in [34, Exercise 12.7] and [6, Chapter 5].
It is proved in [26, Theorem 1] that $F\left(X\right)$ endowed with the Fell topology is a compact Hausdorff space, which is easily seen to be second countable when $X$ is countably compactly based. By identifying a point $x\in X$ with the closed subspace $\overline{\left\{x\right\}}$ of $X$, one can regard $X$ as a subspace of $F\left(X\right)$.
@@
It follows from the above remarks that $\Phi\left(X\right)$ is a compact second countable Hausdorff space. Its elements are characterized in [26, Lemma 2] as the limits sets of Fell-convergent sequences in $X$. Furthermore, a sequence $\left(x_{n}\right)$ in $X$ is Fell-convergent with limit set $C$ if and only if it converges to $C$ in $\Phi\left(X\right)$ [26, Lemma 2].
@@ 4.2. The Fell topology
Notice that the Fell topology on $X$ is in general finer than the Jacobson topology of $X$. Recall that a topological space is Polish if it has a countable basis of open sets and its topology is induced by a complete metric. A subspace of a Polish space is Polish with respect to the subspace topology if and only if it is $G_{\delta}$ [34, Theorem 3.11]. Every locally compact second countable Hausdorff space is Polish [34, Theorem 5.3].
@@
The proof above can be seen as a general version of the proof considered in [17, 1.9.13 and 3.9.2 and 3.9.3] in the case when $X=\mathrm{Prim}\left(A\right)$ for some separable C*-algebra AA; see also [19, Section 2].
@@ 4.3. Universal property
A universal property of the Fell compactification was also isolated by Fell in [26, Theorem 2]. Let us say that a function $f:X\rightarrow Y$ between countably compactly based $T_{0}$ spaces is Fell-continuous if it maps a Fell-convergent sequence $\left(x_{n}\right)$ with limit $C$ to Fell-convergent sequence with limit $f\left(C\right)$. In the case of locally compact second countable Hausdorff spaces, this is equivalent to being proper.
@@ 4.5. Fell and Kuratowski
5. Preliminary spaces
In this section, we introduce the notion of weight function—or, briefly, weight—on a countably compactly based $T_{0}$ space. In terms of this notion, we define preliminary spaces.
@@ 5.1. Weight functions
Suppose that $\left(C_{n}\right)$ is a sequence in $\Phi\left(X\right)$ converging to $C$ and such that $\phi\left(C_{n}\right)=d\in\mathbb{N}$ for every $n\in\mathbb{N}$. Fix $\ell\geq 1$ and $y_{1},\ldots,y_{\ell}\in C$. Then there exist open neighborhoods $U_{1},\ldots,U_{\ell}$ of $y_{1},\ldots,y_{\ell}$, respectively, such that
@@
After passing to subsequences, we can assume that $(y_{n}^{\left(k_{i}\right)})$ is Fell-convergent. By hypothesis, we have
@@
5.2. Preliminary spaces
We now introduce the notion of preliminary space in terms of the notions above.
@@ 5.2. Preliminary spaces
A weighted space is a topological space endowed with a distinguished weight. A preliminary space is a weighted space $\left(X,w\right)$ such that:
@@
    $X$ is countably compactly based and $T_{0}$;
@@
    $w$ is finite.
The preliminary space $\left(X,w\right)$ is $n$-homogeneous when $w$ is the rank function constantly equal to $n$.
@@
If $\left(X,w\right)$ is a liminary space, and $C\subseteq X$ is constructible, then again it is established by induction on the complexity of $C$ that $C$ is a preliminary space when endowed with the induced topology and the restriction of $w$ to $C$.
@@
An $n$-homogeneous preliminary space must necessarily be Polish.
As a preliminary space is, by definition, countably compactly based and $T_{0}$, it is also a standard topological space by Corollary 4.4. We now show that it also admits a Polish decomposition in the sense of Definition 3.2.
@@
Let $\left(X,w\right)$ be a preliminary space. Then $X$ admits a Polish decomposition.
@@
By Lemma 3.6, it suffices to prove that every closed subspace of $X$ admits a Polish open subset. Since a closed subspace of $X$ is still preliminary, it suffices to prove that $X$ has a Polish open subset. For $n\in\mathbb{N}$ define
@@
is a nonempty open subset $X$. Furthermore, $w|_{V}$ is constantly equal to $k$. Thus, $V$ is a $k$-homogeneous preliminary space, and hence Polish. ∎
5.3. Cantor–Bendixon analysis of preliminary spaces
Suppose that $\left(X,w\right)$ is a preliminary space. We define a canonical decomposition of $X$, which can be seen as a sort of Cantor–Bendixon analysis of $X$ [34, Theorem 6.4].
We say that an element of $X$ has order 0 and rank $n<\omega$ if it has a neighborhood $U$ that is $n$-homogeneous, i.e., $w|_{U}$ is constantly equal to $n\in\mathbb{N}$. Define then $X_{n}$ to be the (open) set of elements of order 0 and rank $n$. The derivative $X^{1}$ is the (closed) complement of the union of $X_{n}$ for $n\in\mathbb{N}$.
@@ 5.3. Cantor–Bendixon analysis of preliminary spaces
By [34, Theorem 6.9] there exists $\sigma<\omega_{1}$ such that $X^{\sigma}=\varnothing$. The least such a $\sigma$ is called the order of $X$. When $X$ is compact, $\sigma$ must be a successor ordinal. The set $X^{\alpha}$ is called the $\alpha$-th derivative of $X$. An element $x\in X$ has order $\alpha$ and rank $\omega\alpha+n$ if it belongs to $X_{n}^{\alpha}$, i.e., there exists a neighborhood $U$ of $x$ such that $U\cap X^{\alpha}$ is $n$-homogeneous. The rank of XX\ is the supremum of the ranks of its elements.
@@
Let $X,Y$ be preliminary spaces. Let $f:X\rightarrow Y$ be a continuous function such that $w_{X}\leq w_{Y}\circ f$. Then for every $x\in X$,
@@
$r\left(X\right)\leq r\left(Y\right)\text{.}$
@@ 5.4. Cantor–Bendixon analysis of scattered spaces
The Cantor–Bendixon analysis of a scattered locally compact Polish space [34, Section 6.C] can be seen as a particular instance of the one described above. Indeed, let $X$ be a scattered (equivalently, countable) locally compact Polish space. Recall that one defines its derivative
@@
This turns $X$ into a preliminary space, in such a way that its derivatives coincide with the Cantor–Bendixon derivatives.
@@ 6.1. C*-algebra spectra
We record in the following proposition some properties of spectra of C*-algebras. Suppose that $A$ is a separable C*-algebra. Let $P(A)$ be the pure state space of $A$ [48, Section 4.3]. By [48, Proposition 4.3.2], $P\left(A\right)$ is a Polish space when endowed with the w*-topology. We have a canonical map
@@
(1) If $I$ and $J$ are distinct primitive ideals of $A$, then either $I$ is not contained in $J$ or vice versa. This shows that $X$ is $T_{0}$.
@@
Then we have that $U\left(x\right)$ is open and $K\left(x\right)$ is compact [7, Proposition II.6.5.6]. Let $\mathcal{B}$ be the collection of compact sets $K\left(x\right)$ for $x\in\mathcal{A}$. The proof of [7, Proposition II.6.5.6(ii)] shows that $\mathcal{B}$ witnesses that $X$ is countably compactly based.
(2) This map is continuous and open by [48, Theorem 4.3.3]. Furthermore, by [17, 3.9.2], the corresponding quotient Borel structure on $P\left(A\right)$ is standard, and coincides with the Borel structure generated by the (Polish) Fell topology on $\mathrm{Prim}\left(A\right)$. ∎
@@ 6.2. The Fell compactification of the primitive spectrum
see [17, 1.9.13]. With respect to this correspondence, the Fell topology on $F\left(X\right)$ corresponds to the weakest topology that renders all the functions
@@ 6.3. Unitization
Let $A$ be a nonunital C*-algebra. Its unitization $A^{\dagger}$ [7, Section II.1.2] is defined to be the algebra $A\oplus\mathbb{C}$ with coordinate-wise addition and scalar multiplication, and multiplication induced by the identification
@@
This construction can be seen as a noncommutative analogue of the one-point compactification of a space. Indeed, if $A=C_{0}\left(X\right)$ for some locally compact non compact Polish space, then $A^{\dagger}\cong C\left(X^{\dagger}\right)$ where $X^{\dagger}$ is the one-point compactification of $X$; see [7, II.1.2].
@@
  • full if the closed ideal generated by $a$ is all of $A$ [7, II.5.3.10];
@@
If $A$ is a separable C*-algebra, then $\mathrm{Prim}\left(A\right)$ is compact if and only if $A$ has a strictly full element; see [47, Lemma 2.1].
@@
$\left\|x/J\right\|_{A/J}\leq\liminf_{n}\left\|x/J_{n}\right\|$
@@ 6.4. Spectra of preliminary C*-algebras
We show that the spectrum of a preliminary C*-algebra is indeed a preliminary space.
@@
Then $w$ is a finite weight on $X$, and $\left(X,w\right)$ is a preliminary space.
@@
for every irreducible representation $\pi$ of $A$. Thus, it follows from [4, Theorem 2.4] that $w$ is a weight, which is finite by definition of preliminary C*-algebra. The conclusion now follows from Proposition 6.1. ∎
We consider the spectrum of a preliminary C*-algebra as a preliminary space endowed with the canonical weight function defined in Proposition 6.3.

### v4

@@ 3.2. Hausdorff decompositions
We introduce the notion of standard topological space. Recall the notion of standard Borel space from [43, Section 12.B].
@@
This follows from [43, Lemma 13.2 and Lemma 13.3] considering that $X$ has a countable basis of open sets. ∎
@@
This follows from [43, Theorem 6.9]. ∎
@@ 3.3. Constructible sets and Baire spaces
We consider the following terminology about collections of subsets of a given set; see [33, Section 4].
@@
Let $X$ be a topological space. A subspace $A$ of $X$ is called meager if it is contained in the union of a countable family of closed sets with empty interior. A topological space $X$ is Baire if every nonempty open subset $U$ of $X$ is not meager; see [43, Definition 8.2]. A topological space $X$ is hereditarily Baire if each of its closed subspace of $X$ is Baire, when endowed with the subspace topology. It is easily seen by induction on the complexity of $C$ that, if $X$ is a hereditarily Baire space, and $C\subseteq X$ is constructible, then $C$ is hereditarily Baire.
@@ 3.4. Polish presentations
A space with a standard Polish presentation is a space explicitly presented as the quotient $P/E$ of a Polish space $P$ by an equivalence relation $E$ with the property that:
@@
    the quotient map $P\rightarrow P/E$ is (continuous and) open when $P/E$ is endowed with the quotient topology, i.e., the $E$-saturation of an open subset of $P$ is open;
@@
    the quotient Borel structure on $P/E$ induced by the Borel structure on $P$ is standard, and in particular $E$ is a smooth equivalence relation [43, Exercise 18.20].
A space with a standard Polish presentation has a countable basis of open sets, and it is hereditarily Baire. This follows from the fact that every Polish space is second countable and a Baire space [43, Theorem 8.4], and every closed subspace of a Polish space is $G_{\delta}$, and every $G_{\delta}$ subspace of a Polish space is Polish [43, Theorem 3.11].
@@ 3.5. Countably compactly based spaces
We consider the following property for topological spaces, which simultaneously subsumes second countability and local compactness, and it is equivalent to their conjunction in the Hausdorff case. For a set $Y\subseteq X$ we denote by $\mathrm{int}\left(Y\right)$ its interior.
@@
We say that $X$ has a countable basis of compact open sets when one can take the elements of $\mathcal{B}$ both compact and open.
@@ 4.1. Fell spaces and their compactification
where $K\subseteq X$ is compact and $U$ is open; see also [20, 3.9.2]. In the Hausdorff case, this is also considered in [43, Exercise 12.7] and [7, Chapter 5].
It is proved in [29, Theorem 1] that $F\left(X\right)$ endowed with the Fell topology is a compact Hausdorff space, which is easily seen to be second countable when $X$ is countably compactly based. By identifying a point $x\in X$ with the closed subspace $\overline{\left\{x\right\}}$ of $X$, one can regard $X$ as a subspace of $F\left(X\right)$.
@@
It follows from the above remarks that $\Phi\left(X\right)$ is a compact second countable Hausdorff space. Its elements are characterized in [29, Lemma 2] as the limits sets of Fell-convergent sequences in $X$. Furthermore, a sequence $\left(x_{n}\right)$ in $X$ is Fell-convergent with limit set $C$ if and only if it converges to $C$ in $\Phi\left(X\right)$ [29, Lemma 2].
@@ 4.2. The Fell topology
Notice that the Fell topology on $X$ is in general finer than the Jacobson topology of $X$. Recall that a topological space is Polish if it has a countable basis of open sets and its topology is induced by a complete metric. A subspace of a Polish space is Polish with respect to the subspace topology if and only if it is $G_{\delta}$ [43, Theorem 3.11]. Every locally compact second countable Hausdorff space is Polish [43, Theorem 5.3].
@@
The proof above can be seen as a general version of the proof considered in [20, 1.9.13 and 3.9.2 and 3.9.3] in the case when $X=\mathrm{Prim}\left(A\right)$ for some separable C*-algebra AA; see also [22, Section 2].
@@ 4.3. Universal property
A universal property of the Fell compactification was also isolated by Fell in [29, Theorem 2]. Let us say that a function $f:X\rightarrow Y$ between countably compactly based $T_{0}$ spaces is Fell-continuous if it maps a Fell-convergent sequence $\left(x_{n}\right)$ with limit $C$ to a Fell-convergent sequence with limit $f\left(C\right)$. In the case of locally compact second countable Hausdorff spaces, this is equivalent to being proper.
@@ 4.5. Fell and Kuratowski
5. Weighted spaces
In this section, we introduce the notion of weight function—or, briefly, weight—on a countably compactly based $T_{0}$ space. In terms of this notion, we define weighted spaces.
@@ 5.1. Weight functions
Suppose that $\left(C_{n}\right)$ is a sequence in $\Phi\left(X\right)$ converging to $C$ and such that $w\left(C_{n}\right)=d\in\mathbb{N}$ for every $n\in\mathbb{N}$. Fix $\ell\geq 1$ and $y_{1},\ldots,y_{\ell}\in C$. Then there exist open neighborhoods $U_{1},\ldots,U_{\ell}$ of $y_{1},\ldots,y_{\ell}$, respectively, such that
@@
After passing to subsequences, we can assume that $(y_{n}^{\left(i_{k}\right)})$ is Fell-convergent for every $k\in\left\{1,2,\ldots,m\right\}$. By hypothesis, we have
@@
5.2. Weighted spaces
We now introduce the notion of weighted space in terms of the notions above.
@@ 5.2. Weighted spaces
A weighted space is a pair $\left(X,w\right)$ such that:
@@
    $X$ is countably compactly based and $T_{0}$ topological space;
@@
    $w$ is a finite weight on $X$.
A weighed space $\left(X,w\right)$ is $n$-homogeneous when $w$ is the rank function constantly equal to $n$.
@@
If $\left(X,w\right)$ is a weighted space, and $C\subseteq X$ is constructible, then it is established by induction on the complexity of $C$ that $C$ is a weighted space when endowed with the induced topology and the restriction of $w$ to $C$.
@@
An $n$-homogeneous weighted space must necessarily be Polish.
As a weighted space is, by definition, countably compactly based and $T_{0}$, it is also a standard topological space by Corollary 4.4. We now show that it also admits a Polish decomposition in the sense of Definition 3.2.
@@
Let $\left(X,w\right)$ be a weighted space. Then $X$ admits a Polish decomposition.
@@
By Lemma 3.6, it suffices to prove that every closed subspace of $X$ admits a nonempty Polish open subset. Since a closed subspace of $X$ is still a weighted space, it suffices to prove that $X$ has a nonempty Polish open subset. For $n\in\mathbb{N}$ define
@@
is a nonempty open subset $X$. Furthermore, $w|_{V}$ is constantly equal to $k$. Thus, $V$ is a $k$-homogeneous weighted space, and hence Polish. ∎
5.3. Canonical decomposition of finitely weighted spaces
Suppose that $\left(X,w\right)$ is a weighted space. We define a canonical decomposition of $X$, which can be seen as a sort of Cantor–Bendixon analysis of $X$ [43, Theorem 6.4].
We say that an element of $X$ is locally $n$-homogeneous if it has a neighborhood $U$ that is $n$-homogeneous, i.e., $w|_{U}$ is constantly equal to $n\in\mathbb{N}$. Define then $X_{n}$ to be the (open) set of elements of locally $n$-homogeneous elements. The weighted derivative $X^{1}$ is the (closed) complement of the union of $X_{n}$ for $n\in\mathbb{N}$.
@@ 5.3. Canonical decomposition of finitely weighted spaces
By [43, Theorem 6.9] there exists $\sigma<\omega_{1}$ such that $X^{\sigma}=\varnothing$, which is called the order of $X$. If $\sigma$ is limit, then the rank of $X$ is $\omega\sigma$. If $\sigma$ is successor, then the rank of $X$ is the supremum of $\omega\left(\sigma-1\right)+n$ such that $X_{n}^{\sigma-1}$ is nonempty.
@@
Let $X,Y$ be weighted spaces. Let $f:X\rightarrow Y$ be a continuous function such that $w_{X}\leq w_{Y}\circ f$. Then for every $x\in X$,
@@
$r\left(X,w_{X}\right)\leq r\left(Y,w_{Y}\right)\text{.}$
@@ 5.4. Cantor–Bendixon analysis of scattered spaces
The Cantor–Bendixon analysis of a scattered locally compact Polish space [43, Section 6.C] can be seen as a particular instance of the one described above. Indeed, let $X$ be a scattered (equivalently, countable) locally compact Polish space. Recall that one defines its derivative
@@
This turns $X$ into a weighted space, in such a way that its weighted derivatives coincide with the Cantor–Bendixon derivatives.
@@ 6.1. C*-algebra spectra
We record in the following proposition some properties of spectra of C*-algebras. Suppose that $A$ is a separable C*-algebra. Let $P(A)$ be the pure state space of $A$ [60, Section 4.3]. By [60, Proposition 4.3.2], $P\left(A\right)$ is a Polish space when endowed with the w*-topology. We have a canonical map
@@
If $I$ and $J$ are distinct primitive ideals of $A$, then either $I$ is not contained in $J$ or vice versa. This shows that $X$ is $T_{0}$.
@@
Then we have that $U\left(x\right)$ is open and $K\left(x\right)$ is compact [9, Proposition II.6.5.6]. Let $\mathcal{B}$ be the collection of compact sets $K\left(x\right)$ for $x\in\mathcal{A}$. The proof of [9, Proposition II.6.5.6(ii)] shows that $\mathcal{B}$ witnesses that $X$ is countably compactly based.
This map is continuous and open by [60, Theorem 4.3.3]. Furthermore, by [20, 3.9.2], the corresponding quotient Borel structure on $P\left(A\right)$ is standard, and coincides with the Borel structure generated by the (Polish) Fell topology on $\mathrm{Prim}\left(A\right)$.
By (2), in particular we have that $X$ is hereditarily Baire. ∎
@@ 6.2. The Fell compactification of the primitive spectrum
see [20, 1.9.13]. With respect to this correspondence, the Fell topology on $F\left(X\right)$ corresponds to the weakest topology that renders all the functions
@@ 6.3. Unitization
Let $A$ be a nonunital C*-algebra. Its unitization $A^{\dagger}$ [9, Section II.1.2] is defined to be the algebra $A\oplus\mathbb{C}$ with coordinate-wise addition and scalar multiplication, and multiplication induced by the identification
@@
This construction can be seen as a noncommutative analogue of the one-point compactification of a space. Indeed, if $A=C_{0}\left(X\right)$ for some locally compact non compact Polish space, then $A^{\dagger}\cong C\left(X^{\dagger}\right)$ where $X^{\dagger}$ is the one-point compactification of $X$; see [9, II.1.2].
@@
  • full if the closed ideal generated by $a$ is all of $A$ [9, II.5.3.10];
@@
If $A$ is a separable C*-algebra, then $\mathrm{Prim}\left(A\right)$ is compact if and only if $A$ has a strictly full element; see [59, Lemma 2.1].
@@
$\left\|x/J\right\|\leq\liminf_{n}\left\|x/J_{n}\right\|$
@@ 6.4. Spectra of preliminary C*-algebras
We show that the spectrum of a preliminary C*-algebra is indeed a weighted space.
@@
Then $w$ is a finite weight on $X$, and $\left(X,w\right)$ is a weighted space.
@@
for every irreducible representation $\pi$ of $A$. Thus, it follows from [5, Theorem 2.4] that $w$ is a weight, which is finite by definition of preliminary C*-algebra. The conclusion now follows from Proposition 6.1. ∎
We consider the spectrum of a preliminary C*-algebra as a weighted space endowed with the canonical weight function defined in Proposition 6.3.
Lemma 6.4.
Suppose that $A$ is a separable preliminary C*-algebra, and $J$ is an ideal of $A$. Then:
  1. (1)
    the canonical map from irreducible representations of $A/J$ to irreducible representations of $A$ induces an isomorphism of weighted spaces from $\mathrm{Prim}\left(A/J\right)$ to the closed subspace $\mathrm{hull}\left(J\right)\subseteq\mathrm{Prim}\left(A\right)$;
  2. (2)
    the canonical restriction map from irreducible representations of $A$ that do not vanish on $J$ to irreducible representations of $J$ induces an isomorphism of weighted spaces from the open subspace $\mathrm{Prim}\left(A\right)\setminus\mathrm{hull}\left(A\right)\subseteq\mathrm{Prim}\left(A\right)$ to $\mathrm{Prim}\left(J\right)$.
Proof.
This is an immediate consequence of [9, II.6.1.3 and II.6.1.6]. ∎
6.5. Hausdorff spectrum
[…]
It follows from the Dauns–Hofmann Theorem [13] that for primitive ideals $P,Q$ of $A$, $P\thickapprox Q$ if and only if $P\cap ZA=Q\cap ZA$. In particular, $X$ is Hausdorff if and only if $P\thickapprox Q$ for all $P,Q\in X$, if and only if the function $\mathrm{Prim}\left(A\right)\rightarrow\mathrm{Prim}\left(ZA\right)$, $P\mapsto P\cap ZA$ is injective, in which case it is in fact a homeomorphism $\mathrm{Prim}\left(A\right)\rightarrow\mathrm{Prim}\left(ZA\right)$ [40, Theorem 9.1]. Notice that $P\cap ZA=ZP$ for $P\in X$ by [1, Lemma 3.3].